Insert the indicated number of arithmetic means given the first and last term, show the solution
3 and 36
Indicated 7

Answer by greenestamps(13200)



When n arithmetic means are inserted between a and b, b is equal to a plus (n+1) times the common difference. So in this problem

36 = 3+(7+1)d
36 = 3+8d
8d=33
d=33/8 or 4 1/8

The sequence is

3, 7 1/8, 11 1/4, ..., 31 7/8, 36